### v1

@@ 7. Key estimates for intervals of $\mathcal{\widehat{P}}_{n+1}$
    On the other hand, by Mean Value Theorem, $M$-controlled hypothesis (Theorem 2.6) and the asumption we have
@@
    If $|B_{m,f}(I)|=|B_{m,f}(v)-B_{m,f}(w)|\leq\lambda_{2}^{n/2}$, then by Mean Value Theorem and by exponential converge of renormalizations there exist $a\in B_{m,g}(h(I))$, $b\in B_{m,f}(I)$ and $\beta$ between $a$ and $b$ (note that $\beta$ can not be a critical point of $\mathcal{R}^{m}f$) such that
@@
There exist a universal constant $C_{4}=C_{4}(C,C_{2},C_{3})>0$ and $\lambda_{3}=\lambda_{3}(\lambda_{2})\in(0,1)$ such that for any $I\in\widehat{\mathcal{P}}_{n+1}$, with $I\subset J_{n}$, we have:
@@
By Inequality (LABEL:eq:carlesonpaso1), we just need to prove that for any $I\in\widehat{\mathcal{P}}_{n+1}\setminus\{I_{n+1},I_{n+2}\}$, contained in $I_{n}$, there exist $\widetilde{\lambda}=\widetilde{\lambda}(\lambda_{2},\lambda_{3})\in(0,1)$ and $\widetilde{C}>0$ such that
@@
Let us prove Inequality (11) for $I\in\left\{\{\widehat{\Delta}_{n}^{-j}\}_{j=1}^{j=r(n)-1},\,\Delta_{R}^{n+1},\{\Delta_{1}^{j}\}_{j=0}^{j=r(n)-1}\right\}$, the proof is analogous for intervals in the set $\left\{I_{n+2}(c_{0}),\{\Delta_{a_{n+1}}^{j}\}_{j=0}^{j=\ell(n)-1},\,\Delta_{L}^{n+1},\{\widehat{\Delta}_{n}^{j}\}_{j=0}^{j=\ell(n)}\right\}$. By Proposition 6.1 and the fact that both lenghts $|\widehat{\Delta}_{n}^{-1}|$ and $|\Delta_{1}|$ are comparable with $J_{n}$, we have that both intervals, $\Delta_{1}$ and $\widehat{\Delta}_{n}^{-1}$, satisfy inequality (11) for $\widetilde{C}=2C_{2}C$ and $\widetilde{\lambda}=\lambda_{2}$. Let $M>0$ given by Theorem 2.6 and let $L=L(M)>0$ given by Lemma 5.1. By Proposition 7.1 (with $m=n$ for $I=\Delta_{1}$ and $I=f^{-q_{n+1}}(\widehat{\Delta}_{n}^{-1})$), iterating the map $\mathcal{R}^{n}f$ exactly $L$ times forward from $\Delta_{1}$ and $L$ times backward from $\widehat{\Delta}_{n}^{-1}$, we have for $1\leq j\leq L$,
@@
Otherwise, if $r(n)>2(L+P)$ then the intervals $\Delta_{1}^{j}$ and $\widehat{\Delta}_{n}^{-j}$, for $L+P\leq j$, are all inside the tubular $M_{n,2L}^{f}$ whose center $z_{n}$ is defined. At this point we change of coordinates and consider intervals whose endpoints are given by the tubular-coordinates (see Definition 5.1), i.e. let $\mathcal{F}_{n}=\varphi\circ\mathcal{R}^{n}_{f}\circ\varphi^{-1}$, where $\varphi(x)=D^{2}\mathcal{R}^{n}_{f}(z_{n})(x-z_{n})/2$, and take $|\mathcal{I}_{j}|=|\mathcal{F}_{n}^{j+1}(\varphi(1))-\mathcal{F}_{n}^{j}(\varphi(1))|$ and $|\mathcal{J}_{j}|=|\mathcal{F}_{n}^{-j}(\phi(B_{n,f}(\mathfrak{c}_{n}))-\mathcal{F}_{n}^{-j-1}(\phi(B_{n,f}(\mathfrak{c}_{n}))|$. We also consider $\mathcal{G}_{n}=\psi\circ\mathcal{R}^{n}_{g}\circ\psi^{-1}$, where $\psi(x)=\frac{D^{2}\mathcal{R}^{n}_{g}(h(z_{n}))(x-h(z_{n}))}{2}$.
@@
Using inequality (6) with $s_{0}=\mathcal{F}_{n}^{L+P}(\phi(1))\in(0,1)$, we obtain for $L+P\leq i\leq\min\{i_{r},j_{r}\}$
@@
Therefore, for $L+P\leq i\leq\min\{i_{r},j_{r}\}$
$\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}\right|=\left|\log\frac{|h(s_{i})-h(s_{i+1})|}{|s_{i}-s_{i+1}|}\right|\leq 4(L+1)C_{3}C\lambda_{*}^{n/2}+D_{2}\lambda_{*}^{n/2}$
Analogously, using inequality (6) this time with $s_{0}=\mathcal{F}^{-(L+P)}(\phi(B_{n,f}(\mathfrak{c}_{n}))$), we obtain for any $L+P\leq i\leq\min\{\widehat{i_{\ell}},\widehat{j_{\ell}}\}$
$\left|\log\frac{|h(\mathcal{J}_{i})|}{|\mathcal{J}_{i}|}\right|\leq 4(L+1)C_{3}C\lambda_{*}^{n/2}+D_{2}\lambda_{*}^{n/2}.$
Now, for the intervals inside the Tunnel, and using the calculations in [KT07, Pages 208-209], we obtain for any $\min\{i_{\ell},j_{\ell}\}\leq i\leq\max\{i_{r},j_{r}\}$
$\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}\right|\leq\widetilde{C}_{4}\,\lambda_{*}^{n/2},$
where $\widetilde{C}_{4}=\widetilde{C}_{4}(L,D_{4},C,C_{3})>0$.
Analogously, we obtain the same estimates for $\left|\log\frac{|h(\mathcal{J}_{i})|}{|\mathcal{J}_{i}|}\right|$ for $\min\{\widehat{i_{\ell}},\widehat{j_{\ell}}\}\leq\max\{\widehat{i_{r}},\widehat{j_{r}}\}$. Note also that by the estimatives above and Remark 3.1 we have the proposition also for $I=\Delta_{R}^{n+1}$. We obtain the result for $\lambda_{3}=\max\{\lambda_{2}^{1/2},\lambda^{1/2},\lambda_{*}^{n/2}\}$ and $C_{4}>0$ depending on $L,D_{2},C,C_{3}$ and $D_{4}$. ∎
@@
There exists $C_{5}=C_{5}(M,D_{1},C_{3})>1$ and $\lambda_{5}\in(0,1)$ such that for all $(1-b_{1})n\leq m\leq n-1$, $I\in\widehat{\mathcal{P}}_{n+1}$ and $I\subset J_{m}$ we have:
@@
Note that there are not critical points of $\mathcal{R}^{m}f$ between $\theta$ and $\widetilde{\theta}$ since, in the worse case scenario, $I_{i}$ and $J_{i}$ have $B_{m,f}(\mathfrak{c}_{m})$ as an endpoint in common and then $B_{m,f}(\mathfrak{c}_{m})$ cannot be between $\theta$ and $\widetilde{\theta}$ [^2] Note that if we were using the classical dynamical partitions it could happen that $B_{m,f}(\mathfrak{c}_{m})$ is between $\theta$ and $\widetilde{\theta}$.. Using again Mean Value Theorem there exists $\widehat{\theta}$, between $\theta$ and $\widetilde{\theta}$, such that
@@
Consequentely inequality (20) becomes
@@
Finally, from (22) and (23) we have that inequality 19 becomes
@@
Let us finish this section saying that we can obtain the same results in sections $6$ and $7$ but this time reeplacing the critical point of $f$ $c_{0}$ by $c_{1}$ and $J_{m}(c_{0})$ by $J_{m}(c_{1})$. Moreover, let us mention that it is possible to get the results in sections $6$ and $7$, excepts Lemma 7.2, for the sequence of classical dynamical partition $\{\mathcal{P}_{n}\}$.
@@ 8. Proof of Main Theorem
In this case, $I$ and $I^{\prime}$ are also contained in the same interval of $\widehat{\mathcal{P}}_{n-\lceil b_{2}n\rceil}$, let $\Delta\in\widehat{\mathcal{P}}_{n-\lceil b_{2}n\rceil}$ be such interval. Let $\Delta^{*}$ be the union of $\Delta$ with its left and right neighbours in the partition $\widehat{\mathcal{P}}_{n-\lceil b_{2}n\rceil}$. Let $j<q_{n-\lceil b_{2}n\rceil+1}$ be the smallest natural number such that $f^{j}|_{\Delta^{*}}$ is a diffeomorphism with $f^{j}(\Delta)\subseteq J_{n-\lceil b_{2}n\rceil}(c_{i})$, either for $i=0$ or $i=1$. We observe that $f^{j}(\Delta^{*})$ contains a $\tau$-scaled neighborhood of $f^{j}(I\cup I^{\prime})$: by real bounds and Corollary 3.1, there exist $\widehat{C_{2}}=\widehat{C_{2}}(\widehat{C})>1$ and $\widehat{\mu}\in(0,1)$ such that
@@
and By Koebe Principle there exists $\widehat{C_{3}}=\widehat{C_{3}}(\widehat{C_{2}},\widehat{C})>1$ such that,
@@
then by Koebe distrotion principle and for $x\in I$ and $u\in I^{\prime}$ given by Mean Value Theorem we obtain
@@ 9. Some comments about the multi-critical case
We note that the criterion for smoothness given by Proposition 4.1 is quite general, and therefore it is independent of the quantity of critical points. Let us also recall that Proposition 4.1 is true even if adjacent intervals of the circle partition are not comparable (see also [KT07, Remark 4]). Moreover, the construction of the two-bridges partition can be extend to the multi-critical setting: if the pre-images of the critical points inside $I_{n}(c_{0})$ are all well separed, that means that the exists an interval of the form $[x,f^{q_{n+1}}(x)]$ between them, then the construction is very similar. If some of those pre-images are arbitrarily close to each other (inside or not of the same dynamical interval), we define those pre-images as endpoints of the multi-bridge partition (the endpoints of the classical dynamical interval that contains them will not be endpoints of the multi-bridge partition). In the last case we will have at least one interval in the multi-bridge dynamical partition which has arbitrarily small lenght. In this case, the multi-bridge partition will not satisfy the real bounds property (Theorem 2.1). Despite of that, we can obtain the same estimates given by Proposition 6.1 for endpoints in the multi-bridge partition. However, we do not know if the estimates given by propositions 7.1, 7.2 and Lemma 7.2 are true for arbitrary small length intervals in the multi-bridge partition. Proposition 7.3 is a consequence of the results before.
Acknowledgements

### v2

@@ 7. Key estimates for intervals of $\mathcal{\widehat{P}}_{n+1}$
    On the other hand, by Mean Value Theorem, $M$-controlled hypothesis (Theorem 2.6) and the assumption we have
@@
    If $|B_{m,f}(I)|=|B_{m,f}(v)-B_{m,f}(w)|\leq\lambda_{2}^{n/2}$, then by the Mean Value Theorem and the exponential converge of renormalizations there exist $a\in B_{m,g}(h(I))$, $b\in B_{m,f}(I)$ and $\beta$ between $a$ and $b$ (note that $\beta$ can not be a critical point of $\mathcal{R}^{m}f$) such that
@@
There exist a universal constant $C_{4}=C_{4}(C,C_{2},C_{3})>0$ and $\lambda_{3}=\lambda_{3}(\lambda_{2})\in(0,1)$ such that for any $I\in\widehat{\mathcal{P}}_{n+1}$, with $I\subset J_{n}$,
@@
By Inequality (2), we just need to prove that for any $I\in\widehat{\mathcal{P}}_{n+1}\setminus\{I_{n+1},I_{n+2}\}$, contained in $I_{n}$, there exist $\widetilde{\lambda}=\widetilde{\lambda}(\lambda_{2},\lambda_{3})\in(0,1)$ and $\widetilde{C}>0$ such that
@@
Let us prove Inequality (12) for $I\in\left\{\{\widehat{\Delta}_{n}^{-j}\}_{j=1}^{j=r(n)-1},\,\Delta_{R}^{n+1},\{\Delta_{1}^{j}\}_{j=0}^{j=r(n)-1}\right\}$, the proof is analogous for intervals in the set $\left\{I_{n+2}(c_{0}),\{\Delta_{a_{n+1}}^{j}\}_{j=0}^{j=\ell(n)-1},\,\Delta_{L}^{n+1},\{\widehat{\Delta}_{n}^{j}\}_{j=0}^{j=\ell(n)}\right\}$. By Proposition 6.1 and the fact that both lengths $|\widehat{\Delta}_{n}^{-1}|$ and $|\Delta_{1}|$ are comparable with $J_{n}$, we have that both intervals, $\Delta_{1}$ and $\widehat{\Delta}_{n}^{-1}$, satisfy inequality (12) for $\widetilde{C}=2C_{2}C$ and $\widetilde{\lambda}=\lambda_{2}$. Let $M>0$ given by Theorem 2.6 and let $L=L(M)>0$ given by Lemma 5.1. By Proposition 7.1 (with $m=n$ for $I=\Delta_{1}$ and $I=f^{-q_{n+1}}(\widehat{\Delta}_{n}^{-1})$), iterating the map $\mathcal{R}^{n}f$ exactly $L$ times forward from $\Delta_{1}$ and $L$ times backward from $\widehat{\Delta}_{n}^{-1}$, we have for $1\leq j\leq L$,
@@
Otherwise, if $r(n)>2(L+P)$ then the intervals $\Delta_{1}^{j}$ and $\widehat{\Delta}_{n}^{-j}$, for $L+P\leq j$, are all inside the tubular $M_{n,2L}^{f}$ whose center $z_{n}$ is defined. At this point we change of coordinates and consider intervals whose endpoints are given by the tubular-coordinates (see Definition 5.1), i.e. let $\mathcal{F}_{n}=\varphi\circ\mathcal{R}^{n}_{f}\circ\varphi^{-1}$, where $\varphi(x)=D^{2}\mathcal{R}^{n}_{f}(z_{n})(x-z_{n})/2$, and take $|\mathcal{I}_{j}|=|\mathcal{F}_{n}^{j+1}(\varphi(1))-\mathcal{F}_{n}^{j}(\varphi(1))|$ and $|\mathcal{J}_{j}|=|\mathcal{F}_{n}^{-j}(\phi(B_{n,f}(\mathfrak{c}_{n}))-\mathcal{F}_{n}^{-j-1}(\phi(B_{n,f}(\mathfrak{c}_{n}))|$. We also consider $\mathcal{G}_{n}=\psi\circ\mathcal{R}^{n}_{g}\circ\psi^{-1}$, where $\psi(x)=D^{2}\mathcal{R}^{n}_{g}(h(z_{n}))(x-h(z_{n}))/2$.
@@
Using inequality (7) with $s_{0}=\mathcal{F}_{n}^{L+P}(\phi(1))\in(0,1)$, we obtain for $L+P\leq i\leq\max\{i_{r},j_{r}\}$
@@
Therefore, for $L+P\leq i\leq\max\{i_{r},j_{r}\}$
$\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}\right|=\left|\log\frac{|h(s_{i})-h(s_{i+1})|}{|s_{i}-s_{i+1}|}\right|\leq 4(L+1)C_{3}C\lambda_{*}^{n/2}+D_{2}\lambda_{*}^{n/2}\leq\widetilde{C}_{2}\,\lambda_{*}^{n/2}.$
Analogously, this time for $\mathcal{F}_{n}^{-1}$ and $\mathcal{G}_{n}^{-1}$, we have for $\min\{i_{\ell},j_{\ell}\}\leq i\leq L+P$
$\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}\right|\leq\widetilde{C}_{2}\,\lambda_{*}^{n/2}.$
Moreover, using inequality (7) for $\mathcal{F}_{n}^{i}(\phi(B_{n,f}(\mathfrak{c}_{n})))$ we obtain for any $\min\{\widehat{i_{\ell}},\widehat{j_{\ell}}\}\leq i\leq\max\{\widehat{i_{r}},\widehat{j_{r}}\}$
$\left|\log\frac{|h(\mathcal{J}_{i})|}{|\mathcal{J}_{i}|}\right|\leq\widetilde{C}_{2}\,\lambda_{*}^{n/2}.$
Now, for the intervals inside the Tunnel, we will use Lemma 5.4, with $s_{0}=\mathcal{F}_{n}^{i_{c}}(\phi(1))$ with $s_{k}=\mathcal{F}_{n}^{i_{c}+k}(\phi(1))$, for $0\leq k\leq i_{\ell}-i_{c}$, and also for $s_{0}=\mathcal{G}_{n}^{j_{c}}(\psi(1))$ with $s_{j}=\mathcal{G}_{n}^{j_{c}+j}(\psi(1))$, for $0\leq j\leq j_{\ell}-j_{c}$ By Equation (9) in Lemma 5.4 and Lemma 5.5 we have, for $i_{c}\leq i\leq\min\{i_{\ell},j_{\ell}\}$
$\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}\right|$ $\leq\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}-\log\frac{|\mathcal{G}_{n}^{j_{c}+i-i_{c}+1}(\psi(1))-\mathcal{G}_{n}^{j_{c}+i-i_{c}}(\psi(1)))|}{|\mathcal{F}_{n}^{i+1}(\phi(1))-\mathcal{F}_{n}^{i}(\phi(1))|}\right|+\left|\log\frac{|\mathcal{G}_{n}^{j_{c}+i-i_{c}+1}(\psi(1))-\mathcal{G}_{n}^{j_{c}+i-i_{c}}(\psi(1)))|}{|\mathcal{F}_{n}^{i+1}(\phi(1))-\mathcal{F}_{n}^{i}(\phi(1))|}\right|$
$\leq\left|\sum_{m=i}^{j_{c}-i_{c}+i}\log\frac{|\mathcal{G}_{n}(h(s_{m}))-\mathcal{G}_{n}(h(s_{m+1}))|}{|h(s_{m})-h(s_{m+1})|}\right|+\left|\log\frac{|h(\varepsilon_{n})(1+\delta_{i+j_{c}-i_{c}})(\cos\sqrt{\varepsilon_{n}}(i-i_{c}))^{2}|}{|\varepsilon_{n}(1+\delta_{i})(\cos\sqrt{h(\varepsilon_{n})}(i-i_{c}))^{2}|}\right|$
$\leq|j_{c}-i_{c}|\varepsilon_{n}^{\frac{1}{2+\alpha}}+\left|\log\frac{h(\varepsilon_{n})(1+\delta_{i+jc-i_{c}})}{\varepsilon_{n}(1+\delta_{i})}\right|+\left|2\log\frac{\cos\sqrt{\varepsilon_{n}}(i-i_{c})}{\cos\sqrt{h(\varepsilon_{n})}(i-i_{c})}\right|$
$\leq D_{5}\varepsilon_{n}^{\frac{\alpha(\alpha+1)}{2(2+\alpha)}}+D_{5}\varepsilon_{n}^{\alpha/2}+2D_{4}\,\varepsilon_{n}^{\frac{\alpha(\alpha+1)}{2(2+\alpha)}}+\left|2\log\frac{\cos\sqrt{\varepsilon_{n}}(i_{\ell})}{\cos\sqrt{h(\varepsilon_{n})}(i_{\ell})}\right|$
$\leq D_{6}\,\lambda_{*}^{\frac{n\alpha(\alpha+1)}{4(2+\alpha)}}+\widetilde{C}_{2}\,\lambda_{*}^{n/2}$
$\leq\widetilde{C}_{4}\,\lambda_{*}^{n/2},$
where $D_{6}=(D_{5},M,\alpha)>0$. Analogously we obtain a similar inequality for $\max\{i_{r},j_{r}\}\leq i\leq i_{c}$. Therefore, we have for $\max\{i_{r},j_{r}\}\leq i\leq\min\{i_{\ell},j_{\ell}\}$
$\left|\log\frac{|h(\mathcal{I}_{i})|}{|\mathcal{I}_{i}|}\right|\leq\widetilde{C}_{4}\,\lambda_{*}^{n/2}.$
where $\widetilde{C}_{4}=\widetilde{C}_{4}(L,M,D_{4},D_{5},C,C_{3})>0$.
Analogously, we obtain the same estimates for $\left|\log\frac{|h(\mathcal{J}_{i})|}{|\mathcal{J}_{i}|}\right|$ where $\min\{\widehat{i_{\ell}},\widehat{j_{\ell}}\}\leq i\leq\max\{\widehat{i_{r}},\widehat{j_{r}}\}$. Note also that by the estimatives above and Remark 3.1 we have the proposition also for $I=\Delta_{R}^{n+1}$. Finally, we obtain the result putting all previous inequalities together and choosing $\lambda_{3}=\max\{\lambda_{2}^{1/2},\lambda^{1/2},\lambda_{*}^{n/2}\}$ and $C_{4}>0$ depending on $L,M,D_{2},C,C_{3},D_{4}$ and $D_{5}$. ∎
@@
There exists $C_{5}=C_{5}(M,D_{1},C_{3})>1$ and $\lambda_{5}\in(0,1)$ such that for all $(1-b_{1})n\leq m\leq n-1$, $I\in\widehat{\mathcal{P}}_{n+1}$ and $I\subset J_{m}$
@@
Note that there are no critical points of $\mathcal{R}^{m}f$ between $\theta$ and $\widetilde{\theta}$ since, in the worse case scenario, $I_{i}$ and $J_{i}$ have $B_{m,f}(\mathfrak{c}_{m})$ as an endpoint in common and then $B_{m,f}(\mathfrak{c}_{m})$ cannot be between $\theta$ and $\widetilde{\theta}$ [^2] Note that if we were using the classical dynamical partitions it could happen that $B_{m,f}(\mathfrak{c}_{m})$ is between $\theta$ and $\widetilde{\theta}$.. Using the Mean Value Theorem again there exists $\widehat{\theta}$, between $\theta$ and $\widetilde{\theta}$, such that
@@
Consequently inequality (21) becomes
@@
Finally, from (23) and (24) we have that inequality 20 becomes
@@
Let us finish this section saying that we can obtain the same results in sections $6$ and $7$ replacing the critical point $c_{0}$ by $c_{1}$ and $J_{m}(c_{0})$ by $J_{m}(c_{1})$. Moreover, let us mention that it is also possible to get the results in sections $6$ and $7$ for the sequence of classical dynamical partition $\{\mathcal{P}_{n}\}$, excepts Lemma 7.2.
@@ 8. Proof of Main Theorem
In this case, $I$ and $I^{\prime}$ are also contained in the same interval of $\widehat{\mathcal{P}}_{n-\lceil b_{2}n\rceil}$, let $\Delta\in\widehat{\mathcal{P}}_{n-\lceil b_{2}n\rceil}$ be such interval. Let $\Delta^{*}$ be the union of $\Delta$ with its left and right neighbors in the partition $\widehat{\mathcal{P}}_{n-\lceil b_{2}n\rceil}$. Let $j<q_{n-\lceil b_{2}n\rceil+1}$ be the smallest natural number such that $f^{j}|_{\Delta^{*}}$ is a diffeomorphism with $f^{j}(\Delta)\subseteq J_{n-\lceil b_{2}n\rceil}(c_{i})$, either for $i=0$ or $i=1$. We observe that $f^{j}(\Delta^{*})$ contains a $\tau$-scaled neighborhood of $f^{j}(I\cup I^{\prime})$: by real bounds and Corollary 3.1, there exist $\widehat{C_{2}}=\widehat{C_{2}}(\widehat{C})>1$ and $\widehat{\mu}\in(0,1)$ such that
@@
and by the Koebe Principle there exists $\widehat{C_{3}}=\widehat{C_{3}}(\widehat{C_{2}},\widehat{C})>1$ such that,
@@
then by Koebe distortion principle and for $x\in I$ and $u\in I^{\prime}$ given by Mean Value Theorem we obtain
@@ 9. Some comments about the multi-critical case
We note that the criterion for smoothness given by Proposition 4.1 is quite general, and therefore it is independent of the quantity of critical points. Let us also recall that Proposition 4.1 is true even if adjacent intervals of the circle partition are not comparable (see also [KT07, Remark 4]). Moreover, the construction of the two-bridges partition can be extend to the multi-critical setting: if the pre-images of the critical points inside $I_{n}(c_{0})$ are all well separated, that means that the exists an interval of the form $[x,f^{q_{n+1}}(x)]$ between them, then the construction is very similar. If some of those pre-images are arbitrarily close to each other (inside or not of the same dynamical interval), we define those pre-images as endpoints of the multi-bridge partition (the endpoints of the classical dynamical interval that contains them will not be endpoints of the multi-bridge partition). In the last case we will have at least one interval in the multi-bridge dynamical partition which has arbitrarily small length. In this case, the multi-bridge partition will not satisfy the real bounds property (Theorem 2.1). Despite of that, we can obtain the same estimates given by Proposition 6.1 for endpoints in the multi-bridge partition. However, we do not know if the estimates given by propositions 7.1, 7.2 and Lemma 7.2 are true for intervals of arbitrarily small length in the multi-bridge partition. Proposition 7.3 is a consequence of the results before.
Acknowledgments